# Restyling Unsupervised Concept Based Interpretable Networks with Generative Models

### v1

@@ Abstract
Developing inherently interpretable models for prediction has gained prominence in recent years. A subclass of these models, wherein the interpretable network relies on learning high-level concepts, are valued because of closeness of concept representations to human communication. However, the visualization and understanding of the learnt unsupervised dictionary of concepts encounters major limitations, specially for large-scale images. We propose here a novel method that relies on mapping the concept features to the latent space of a pretrained generative model. The use of a generative model enables high quality visualization, and naturally lays out an intuitive and interactive procedure for better interpretation of the learnt concepts. Furthermore, leveraging pretrained generative models has the additional advantage of making the training of the system more efficient. We quantitatively ascertain the efficacy of our method in terms of accuracy of the interpretable prediction network, fidelity of reconstruction, as well as faithfulness and consistency of learnt concepts. The experiments are conducted on multiple image recognition benchmarks for large-scale images. Project page available at https://jayneelparekh.github.io/VisCoIN_project_page/
Keywords:
Interpretability Explainability Generative models
@@ 1 Introduction
Deep neural networks (DNNs) learn complex patterns from data to make predictions or decisions without being explicitly programmed how to perform the task. Interpreting decisions of DNNs, i.e. being able to obtain human-understandable insights about their decisions, is a difficult task [13, 11, 49]. This lack of transparency impacts their trustworthiness [58] and hinders their democratization for critical applications such as assisting medical diagnosis or autonomous driving.
@@
Two different paths have been explored in order to interpret DNNs outputs. The simplest approach for practitioners is to provide interpretations post-hoc, i.e. by analysing the so-called black-box model after training [12, 55, 47, 61]. However, post-hoc methods have been criticized for their high computational costs and a lack of robustness and faithfulness of interpretations [78, 37, 8]. On the other hand, one preferred way to obtain more meaningful interpretations is to use interpretable by-design approaches [5, 3, 14, 21], that aim to integrate the interpretability constraint into the learning process, while maintaining state-of-the-art performance.
@@
Concept-based Interpretable Networks (CoINs) are a recent subcategory of these inherently interpretable prediction models, that learn a dictionary of high-level concepts for prediction. The concept representation is either learnt in a supervised way using ground-truth concepts annotations [40, 59], or in an unsupervised fashion by enforcing properties through carefully designed loss functions [7, 51, 59]. The output of the model can be interpreted by looking at the activations of each concept and how they are combined to obtain the final prediction. When working in the unsupervised setting, learnt concepts have to additionally be interpreted, usually through visualization [51, 59]. Concept-based interpretations have gained prominence as an alternative to popular feature-wise saliency maps [66, 55, 47, 61] for two main reasons: (1) their ability to provide interpretations closer to human reasoning and communication [76], and (2) in specific case of visual modalities, their ability to more effectively highlight which features are important for a model and not just where in the input image they focus on [16]. However, the underlying concepts in current unsupervised CoINs are understood through a separate visualization pipeline, by finding inputs that highly activate a given concept, either from natural images in the available dataset [7, 59], or from virtual images by solving an optimization problem in the input space that maximally activates the concept [48, 51]. For large-scale images, concepts generally activate for local pattern information (color, texture, shape etc.) and these visualization approaches face major limitations in highlighting this information to a user. Simply visualizing the most activating samples does not highlight the specific feature a concept activates for. Visualizing using an activation maximization procedure leads to the generation of repeated patterns linked to the underlying concept in the image, but are hard for a user to discern any human-interpretable signal. For example, in Fig. 1, it can be hard to identify that the concept activates for “Yellow-colored head” from the activation maximization (“FLINT visualization”). Furthermore, previous CoIN systems fail to include the visualization process in their quantitative evaluation of concepts.
@@
We thus propose a novel set of specifications for the concepts to be learnt: additionally to fidelity to output (predictive capability from the concepts), fidelity to input (encoding input relevant information in concepts) and sparsity (a few concepts activated simultaneously), we also promote the viewability of concepts during training. This viewability is now defined as the ability of the system to reconstruct high-quality images from the learnt concepts, by leveraging a pretrained generative model. In order to obtain this viewability property, we propose to learn a concept translator, i.e., a mapping from the concept representation space to the latent space of the generative model. Learning the concept translator along with the other parameters of this novel CoIN system helps to improve the quality of the concepts. Finally, after training, interpretation of concepts is obtained through translation to the generative model, which allows for a more granular and interactive process. Our contributions are the following:
@@
    We introduce new metrics in the context of unsupervised CoINs to evaluate the quality of concepts learnt, from the point of view of visualization. We then quantitatively and qualitatively evaluate our proposed method on three different large-scale image datasets, spanning multiple settings.
Figure 1: Comparison of the generated images obtain for the same learnt concept (“Yellow-colored head”) using FLINT visualization [51] and our proposed VisCoIN visualization. Using our concept translator, that maps concept representation space to the latent space of a generative model, we can visualize each concept at different activation values, allowing for more granular and interactive interpretation.
@@ 2 Related works
Interpretable predictive models . In the context of deep learning architectures, a host of early approaches studying interpretability tackled the post-hoc interpretation problem [63, 66, 55, 47, 67, 15]. However, previous works such as those of [5, 45, 7] have contributed to surge of developing predictive models that are also interpretable by-design [77, 4, 81, 44, 21]. The earlier systems however trained the complete model from scratch. Recent approaches reflect a growing interest in building interpretable models on top of pretrained models as backbones [40, 9]. Our approach falls in the latter category wherein we learn an interpretable predictive model on top of a pretrained backbone.
GANs for interpretations . One of the earliest application of GANs for interpretability was by [50] to synthesize image for visualizing neurons in a network. More recently, a variety of methods have employed GAN or other generative models (including diffusion models) for generating post-hoc counterfactual interpretations [80, 43, 19, 22]. Their central theme revolves around the idea of embedding any given input to the latent space of a GAN and finding meaningful perturbations in the latent space that affect the given predictor’s output the most. Our aimed use-case of the GAN differs in a major way from these methods, which is that we wish to use the GAN in order to learn and visualize an explicit dictionary of interpretable concept representation simultaneously used in a predictive model.
Concept-based interpretability . Providing interpretations via representations for high-level concepts has gained significant prominence recently. Similar to the overall literature, one set of concept-based methods have focused on post-hoc interpretation [23, 76, 43, 2, 20], with most based on the notion of concept activation vectors [36]. The other type of methods tackle the by-design/ante-hoc interpretation problem by learning concepts [7, 40, 51, 59, 60, 62] abbreviated as CoIN systems in Sec. 1. We cover these methods in more detail in Sec. 3.1 with particular focus on networks based on learning completely unsupervised concepts [7, 51, 59], a key starting point of our approach.
@@ 3.1 Background
In this part, we provide a design overview of a concept based interpretable network (CoIN). Our focus in this paper is on CoIN systems that learn an unsupervised dictionary of concepts.
Concept-based interpretable networks. We denote a training set for a supervised image classification task as $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$. Each input image $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is associated with a class label $y\in\mathcal{Y}$, a one-hot vector of size number of classes $C$. The by-design interpretable network based on learning concept representation is denoted by $g:\mathcal{X}\rightarrow\mathcal{Y}$.
In the standard setup for concept-based prediction models (supervised or unsupervised), given an input $x$, the computation of $g(x)$ is broken down into two parts. There is first a concept extraction representation $\Phi$, and then a subnetwork $\Theta$ that computes the final prediction using concept activations $\Phi(x)$, such that $g(x)=\Theta\circ\Phi(x)$. Supervised concept-based networks [40] use ground-truth concept annotations to train $\Phi$. The core of unsupervised concept-based methods instead lies in learning $\Phi$ by imposing loss functions. These loss functions are typically selected to encourage a certain set of properties that shape $\Phi$ simultaneously for both interpretation and prediction. We list the properties below:
@@
    Fidelity to output: This requires $\Phi(x)$ to model the output space via the $\Theta$ function, either by predicting ground-truth label $y$ [7] or the classification output $f(x)$ [51, 59]. It directly shapes $\Phi(x)$ for the prediction task and, during the interpretation phase, it helps in identifying important concepts for prediction.
@@
    Fidelity to input: This requires $\Phi(x)$ to reconstruct the input $x$ via a decoder function. This property is essential to encode features relevant to the input in the concept representation. All the previous methods [7, 51, 59] rely on this loss and employ standard non-generative decoders for pixel-wise reconstruction to learn the concept dictionary $\Phi$.
@@
    Sparsity of activations: This requires concept activations $\Phi(x)$ to be sparse for any $x$. It reinforces the high-level nature of $\Phi$ and limits the number of important concepts for prediction, in turn enhancing interpretability as well as reducing the visualization overhead for a user.
Limitations with concept visualization in previous unsupervised CoINs . A central common trait among prior CoINs learning unsupervised concepts [7, 51, 59] is the deployment of a decoder to reconstruct input $x$ from $\Phi(x)$. Unlike supervised methods, they do not have access to any concept labels and thus need an additional visualization pipeline to understand the information encoded by each concept. However, their visualization pipeline does not utilize the decoder, but instead relies on proxy methods to probe the concept activation. Typically, it consists of finding an input that highly activates a concept, either by selecting from the training data [7, 59] or via input optimization [51]. In the former case, simply visualizing the set of most activating training samples lacks granularity to highlight the features encoded by the concept. Using input optimization, while relatively more insightful, is still difficult for a user to understand as the optimized images are often unnatural. Moreover, these issues exacerbate for large-scale images, as seen in Fig. 1. A natural strategy to overcome these limitations is to enable direct control of a concept’s activation and visualizing its effect on the input. Since the decoder defines the relationship between concept activations and input samples, a generative model is a perfect candidate for a decoder to unlock this ability, in contrast to standard decoders used previously.
In the next part, we describe the architecture behind our by-design interpretable network $g$, that additionally includes a concept translator module $\Omega$, to map concept features in the latent space of a pretrained generative model $G$. This defines our VisCoIN method.
@@ 3.2 Interpretable prediction network design
We assume a fixed pretrained network for classification $f$ and a fixed pretrained generator $G$, trained for classification and generation on input dataset respectively. We use these two networks to guide our design and learning of $g$ and its concept extraction function $\Phi$. We first discuss modelling of $g$ by describing its constituents, $\Phi$ and $\Theta$.
@@
Viewability property . In order to improve visualization for unsupervised CoINs, and thus interpretation of learnt concepts, we propose to add the requirement for a viewability property. Given an input image $x$, this property requires to be able to reconstruct high-quality images from $\Phi(x)$, from a pretrained generative model $G$. More specifically, reconstructions should have high enough quality to “view” input samples through generated outputs and thus ground modifications to $\Phi(x)$ back to $x$. In our method, we propose to achieve this by learning an additional concept translator module $\Omega$ to map $\Phi(x)$ to the latent space of $G$, such that high-quality reconstructed images can be obtained from $\Phi(x)$ through $\Omega$ and $G$. The design of $\Omega$ depends entirely on the generative model used.
We opt for StyleGAN2-ADA [31] as our generative model. Our rationale behind this choice is due to certain properties exhibited by this model: (a) a low dimensional and highly structured latent space with meaningful transformations in generated output via latent traversals, (b) can generate high-quality large-scale images, and (c) flexible to train in limited data regimes. We discuss about the significance of these properties and suitability of other generative models in much greater detail in Sec. 0.A.1. It’s worth noting that using a pretrained generator also limits the training cost, complexity and improves reusability.
3.3 Use of StyleGAN as decoder
A short primer on style-based GAN . The work by Karras et al. [33] proposed a style-based generator architecture. The generator learns to map a noise vector ($z\in\mathbb{R}^{512}\sim\mathcal{N}(0,1)$) to a latent code ($w\in\mathcal{W}\subset\mathbb{R}^{512}$) that controls the scaling and biasing of feature maps at different resolutions in the synthesis network, controlling the style of the generated image. Subsequent updates on the original architecture [34, 32] have improved upon various aspects of generator architecture and regularization. One such update relevant for us is the StyleGAN2-ADA [31] which incorporates an adaptive discriminator augmentation in training. This enables training StyleGAN2 with limited amount of data, a regime where GANs are generally prone to collapsing. The StyleGAN architecture has been actively studied for certain aspects relevant to our use case. Primary among these are GAN inversion methods that aim to embed any given input to the latent space of a pretrained GAN so that the GAN can reconstruct the input using the latent code. The two approaches for this include either optimizing a latent code for any given input [1, 29], or training an encoder to predict the latent code [56, 6, 75]. While the optimization approaches can invert a given input better, they scale poorly compared to encoder based approaches when inverting many input samples. StyleGAN has also been studied closely in regard to its structure of latent spaces, including their suitability to embed inputs [1, 72, 73], discovering semantically meaningful directions or trajectories in latent space [25, 69, 65] and hierarchical style control with latent vectors for different resolutions [35].
At this juncture, it is worth asking the question of why don’t we directly use the latent space/stylespace [72] of the pretrained StyleGAN as the concept dictionary for interpretation, as done previously [43]. We discuss about this question in detail in Sec. 0.A.2, including issues in this design for use in a by-design interpretable network. Instead, we adhere to standard setup for unsupervised CoINs to learn a small dictionary of concept functions. The pretrained StyleGAN is used to reconstruct input $x$ from concept activations $\Phi(x)$.
The most natural way to design this reconstruction is to use a similar approach to encoder based GAN-inversion architectures used for StyleGAN. Following previous works in this regard [1, 75], we use the extended latent space $\mathcal{W}^{+}$ for inversion, which corresponds to different latent vectors for different resolutions. We learn the concept translator $\Omega$ to map the concept activations to $\mathcal{W}^{+}$ and bias its computation with average latent vector $\bar{w}$ of $G$.
Since $\Phi(x)$ is a much lower dimensional representation compared to elements in $\mathcal{W}^{+}$ and constrained by losses unrelated to reconstruction, we found it challenging to achieve reconstruction quality close to GAN inversion methods. This issue in principle cannot be completely eliminated without compromising the interpretable predictive structure. However, we alleviate it to some extent by learning an unconstrained “supporting” representation as a secondary output from $\Psi$, termed $\Phi^{\prime}(x)$. The only goal for $\Phi^{\prime}(x)$ is to assist $\Omega$ in embedding the input in $\mathcal{W}^{+}$. We utilize findings from the work in [35] about hierarchical nature of latent vectors and control prediction of small number of latent vectors using $\Phi^{\prime}(x)$ which are expected to contain minimal classification related information. To ensure importance of $\Phi(x)$ in reconstruction is not diminished, majority of latent vectors are still predicted using $\Phi(x)$. The choice of using $\Phi^{\prime}$ is analyzed with an ablation study. The computation for reconstructed input $\tilde{x}$ is given as:
The concept translator $\Omega$ consists of a set of single fully-connected (FC) layers, one for predicting each latent vector. Each FC layer either takes $\Phi(x)$ or $\Phi^{\prime}(x)$ as input depending upon the latent vector it predicts. Precise details about $\Omega$ and $\Phi^{\prime}(x)$ are in Appendix 0.B. Fig. 2 illustrates the complete design of VisCoIN.

### v2

@@ Abstract
Developing inherently interpretable models for prediction has gained prominence in recent years. A subclass of these models, wherein the interpretable network relies on learning high-level concepts, are valued because of closeness of concept representations to human communication. However, the visualization and understanding of the learnt unsupervised dictionary of concepts encounters major limitations, especially for large-scale images. We propose here a novel method that relies on mapping the concept features to the latent space of a pretrained generative model. The use of a generative model enables high quality visualization, and lays out an intuitive and interactive procedure for better interpretation of the learnt concepts by imputing concept activations and visualizing generated modifications. Furthermore, leveraging pretrained generative models has the additional advantage of making the training of the system more efficient. We quantitatively ascertain the efficacy of our method in terms of accuracy of the interpretable prediction network, fidelity of reconstruction, as well as faithfulness and consistency of learnt concepts. The experiments are conducted on multiple image recognition benchmarks for large-scale images. Project page available at https://jayneelparekh.github.io/VisCoIN_project_page/
@@ 1 Introduction
Deep neural networks (DNNs) learn complex patterns from data to make predictions or decisions without being explicitly programmed how to perform the task. Interpreting decisions of DNNs, i.e. being able to obtain human-understandable insights about their decisions, is a difficult task (Beaudouin et al., 2020; Arrieta et al., 2020; Montavon et al., 2019). This lack of transparency impacts their trustworthiness (Rudin et al., 2022) and hinders their democratization for critical applications such as assisting medical diagnosis or autonomous driving. Two different paths have been explored in order to interpret DNNs outputs. The simplest approach for practitioners is to provide interpretations post-hoc, i.e. by analysing the so-called black-box model after training (Baehrens et al., 2010; Ribeiro et al., 2016; Lundberg & Lee, 2017; Selvaraju et al., 2017). However, post-hoc methods have been criticized for their high computational costs and a lack of robustness and faithfulness of interpretations (Yosinski et al., 2015; Kindermans et al., 2019; Alvarez-Melis & Jaakkola, 2018b). On the other hand, one preferred way to obtain more meaningful interpretations is to use interpretable by-design approaches (Al-Shedivat et al., 2017; Adel et al., 2018; Böhle et al., 2022; Gautam et al., 2022), that aim to integrate the interpretability constraint into the learning process, while maintaining state-of-the-art performance.
@@
Concept-based Interpretable Networks (CoINs) are a recent subcategory of these inherently interpretable prediction models, that learn a dictionary of high-level concepts for prediction. The concept representation is either learnt in a supervised way using ground-truth concepts annotations (Koh et al., 2020; Sarkar et al., 2022), or in an unsupervised fashion by enforcing properties through carefully designed loss functions (Alvarez-Melis & Jaakkola, 2018a; Parekh et al., 2021; Sarkar et al., 2022). The output of the model can be interpreted by looking at the activations of each concept and how they are combined to obtain the final prediction. When working in the unsupervised setting, learnt concepts have to additionally be interpreted, usually through visualization (Parekh et al., 2021; Sarkar et al., 2022). Concept-based interpretations have gained prominence as an alternative to popular feature-wise saliency maps (Springenberg et al., 2014; Ribeiro et al., 2016; Lundberg & Lee, 2017; Selvaraju et al., 2017) for two main reasons: (1) their ability to provide interpretations closer to human reasoning and communication (Yeh et al., 2019), and (2) in specific case of visual modalities, their ability to more effectively highlight which features are important for a model and not just where in the input image they focus on (Colin et al., 2022). However, the underlying concepts in current unsupervised CoINs are understood through a separate visualization pipeline, by finding inputs that highly activate a given concept, either from natural images in the available dataset (Alvarez-Melis & Jaakkola, 2018a; Sarkar et al., 2022), or from virtual images by solving an optimization problem in the input space that maximally activates the concept (Mahendran & Vedaldi, 2016; Parekh et al., 2021). For large-scale images, concepts generally activate for local pattern information (color, texture, shape etc.) and these visualization approaches face major limitations in highlighting this information to a user. Simply visualizing the most activating samples does not highlight the specific feature a concept activates for. Visualizing using an activation maximization procedure leads to the generation of repeated patterns linked to the underlying concept in the image, but are hard for a user to discern any human-interpretable signal. For example, in Fig. 1, it can be hard to identify that the concept activates for “Yellow-colored head” from the activation maximization (“FLINT visualization”). Furthermore, previous CoIN systems fail to include the visualization process in their quantitative evaluation of concepts and their use-case for interpretation.
@@
We thus propose a novel set of specifications for the concepts to be learnt: additionally to fidelity to output (predictive capability from the concepts), fidelity to input (encoding input relevant information in concepts) and sparsity (a few concepts activated simultaneously), we also promote the viewability of concepts during training. This viewability is now defined as the ability of the system to reconstruct high-quality images from the learnt concepts, by leveraging a pretrained generative model. In order to obtain this viewability property, we propose to learn a concept translator, i.e., a mapping from the concept representation space to the latent space of the generative model. Learning the concept translator along with the other parameters of this novel CoIN system helps to improve the quality of the concepts. Finally, interpretation of concepts is obtained through translation to the generative model, allowing for a more granular and interactive process. Our contributions are:
@@
    We introduce new metrics in the context of unsupervised CoINs to evaluate the quality of concepts learnt, from the point of view of visualization and its usage for interpretation. We then quantitatively and qualitatively evaluate our proposed method on three different large-scale image datasets, spanning multiple settings.
Figure 1: Comparison of the generated images obtain for the same learnt concept (“Yellow-colored head”) using FLINT visualization (Parekh et al., 2021) and our proposed VisCoIN visualization (in red boxes). Using our concept translator, that maps concept representation space to the latent space of a generative model, we can visualize each concept at different activation values, allowing for more granular and interactive interpretation. Visual modifications manually indicated by red boxes.
@@ 2 Related works
Interpretable predictive models . In the context of deep learning architectures, a host of early approaches studying interpretability tackled the post-hoc interpretation problem (Simonyan et al., 2013; Springenberg et al., 2014; Ribeiro et al., 2016; Lundberg & Lee, 2017; Sundararajan et al., 2017; Chen et al., 2018). However, previous works such as those of Al-Shedivat et al. (2017); Li et al. (2018); Alvarez-Melis & Jaakkola (2018a) have contributed to surge of developing predictive models that are also interpretable by-design (Yoon et al., 2018; Agarwal et al., 2020; Zhang et al., 2018a; Lee et al., 2019; Gautam et al., 2022). The earlier systems, however, trained the complete model from scratch. Recent approaches reflect a growing interest in building interpretable models on top of pretrained models (Koh et al., 2020; Angelov et al., 2023). Our approach falls in the latter category, wherein we learn an interpretable predictive model on top of a pretrained backbone.
Generative models for interpretations . One of the earliest applications of generative models for interpretability was by Nguyen et al. (2016) to synthesize image for visualizing neurons in a network using GANs. More recently, a variety of methods have employed generative models for post-hoc counterfactual interpretations (Zemni et al., 2022; Lang et al., 2021; Farid et al., 2023; Ghandeharioun et al., 2021). Their central theme revolves around the idea of embedding any given input to the latent space of a generative model and finding meaningful perturbations in the latent space that affect the given predictor’s output the most. One recent work (Ismail et al., 2023) also included the task of learning supervised concepts within generative models, to be able to interpret and steer their latent spaces. Our aimed use-case of generative models differs in a major way from these methods, because we wish to use it in order to learn and visualize an explicit dictionary of interpretable concept representation, simultaneously used in a predictive model.
Concept-based interpretability . Providing interpretations via representations of high-level concepts has gained significant prominence recently. Similar to the overall literature, one set of concept-based methods have focused on post-hoc interpretation (Ghorbani et al., 2019; Yeh et al., 2019; Lang et al., 2021; Achtibat et al., 2022; Fel et al., 2023), with most based on the notion of concept activation vectors (Kim et al., 2017). The other type of methods tackle the by-design/ante-hoc interpretation problem by learning concepts (Alvarez-Melis & Jaakkola, 2018a; Koh et al., 2020; Parekh et al., 2021; Sarkar et al., 2022; Sawada & Nakamura, 2022; Sheth & Ebrahimi Kahou, 2024) abbreviated as CoIN systems in Section 1. We cover these methods in more detail in Section 3.1 with particular focus on networks based on learning completely unsupervised concepts (Alvarez-Melis & Jaakkola, 2018a; Parekh et al., 2021; Sarkar et al., 2022; Garg et al., 2024), a key starting point of our approach. Recent variants of concept bottleneck models using language models (Oikarinen et al., 2023; Panousis et al., 2024) and why concept visualization is still useful despite their ability to obtain automated text descriptions, is discussed in detail in Appendix A.
@@ 3.1 Background
In this part, we provide an overview of a concept based interpretable network (CoIN). Our focus in this paper is on CoIN systems that learn an unsupervised dictionary of concepts.
Concept-based interpretable networks. We denote a training set for a supervised image classification task as $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$. Each input image $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is associated with a class label $y\in\mathcal{Y}$, a one-hot vector of size number of classes $C$. The by-design interpretable classification network based on learning concept representation is denoted by $g:\mathcal{X}\rightarrow\mathcal{Y}$. In the standard setup for concept-based prediction models (supervised or unsupervised), given an input $x$, the computation of $g(x)$ is broken down into two parts. There is first a concept extraction representation $\Phi$, and then a subnetwork $\Theta$ that computes the final prediction using concept activations $\Phi(x)$, such that $g(x)=\Theta\circ\Phi(x)$ (Fig. 2, left). Supervised concept-based networks use sample-wise ground-truth concept annotations to train $\Phi$. For instance, concept bottleneck models (CBM) (Koh et al., 2020) use a human-annotated binary concept vector denoting presence/absence of each concept. Although language model based CBMs do not require manual annotation of concepts, they also follow a very similar paradigm, e.g. by using CLIP “image-concept description” similarities as proxy annotations (Oikarinen et al., 2023; Yang et al., 2023a; Panousis et al., 2024). The core of unsupervised concept-based methods instead lies in learning $\Phi$ by imposing loss functions. These loss functions are typically selected to encourage a certain set of properties that shape $\Phi$ simultaneously for both interpretation and prediction. We list the properties below:
@@
    Fidelity to output: This requires $\Phi(x)$ to model the output space via the $\Theta$ function, either by predicting ground-truth label $y$ (Alvarez-Melis & Jaakkola, 2018a) or the classification output $f(x)$ (Parekh et al., 2021; Sarkar et al., 2022). It trains $\Phi(x)$ for the prediction task and, during the interpretation phase, helps in identifying important concepts for prediction.
@@
    Fidelity to input: This requires $\Phi(x)$ to reconstruct the input $x$ via a decoder function. This property is considered important to encode semantically meaningful input features in the concept representation. All the previous methods (Alvarez-Melis & Jaakkola, 2018a; Parekh et al., 2021; Sarkar et al., 2022) rely on this loss and employ standard non-generative decoders for pixel-wise reconstruction to learn the concept dictionary $\Phi$. Note that CBMs don’t adhere to this property, leading to major design and training differences, as they don’t include a decoder compared to unsupervised CoINs, that remain the focus of this work.
@@
    Sparsity of activations: This requires concept activations $\Phi(x)$ to be sparse for any $x$. It reinforces the high-level nature of $\Phi$ and limits the number of important concepts for prediction, thus enhancing interpretability and reducing the visualization overhead for users.
Limitations with concept visualization in previous unsupervised CoINs . A common trait among prior CoINs learning unsupervised concepts (Alvarez-Melis & Jaakkola, 2018a; Parekh et al., 2021; Sarkar et al., 2022; Garg et al., 2024) is the deployment of a decoder to reconstruct input $x$ from $\Phi(x)$. Unlike supervised methods, they do not have access to any concept labels and thus need an additional visualization pipeline to understand the information encoded by each concept. However, their visualization pipeline does not utilize the decoder, but instead relies on proxy methods to probe the concept activation. Typically, it consists of finding an input that highly activates a concept, either by selecting from the training data (Alvarez-Melis & Jaakkola, 2018a; Sarkar et al., 2022) or via input optimization (Parekh et al., 2021). In the former case, simply visualizing the set of most activating training samples lacks granularity to highlight the features encoded by the concept. Using input optimization, while relatively more insightful, is still difficult for a user to understand as the optimized images are often unnatural. Moreover, these issues exacerbate for large-scale images, as seen in Fig. 1. A natural strategy to overcome these limitations is to enable direct control of a concept’s activation and visualizing its effect on the input. Since the decoder defines the relationship between concept activations and input samples, a generative model is a perfect candidate for a decoder to unlock this ability, in contrast to standard decoders used previously. While Garg et al. (2024) includes a GAN as a decoder model, it is learned simultaneously. This makes the overall training challenging, since GANs are notoriously difficult to train. Furthermore, they do not leverage the GAN for visualizing the concepts, only relying on maximum activating samples (MAS) for visualization. The GAN is used as a decoder with higher expressivity, to improve accuracy.
In the next part, we describe the architecture behind our by-design interpretable network $g$, that additionally includes a concept translator module $\Omega$, to map concept features $\Phi(x)$ to the latent space of a pretrained generative model $G$. This collectively defines our VisCoIN method.
@@ 3.2 Interpretable prediction network design
The complete design of VisCoIN is illustrated in Fig. 2 (right). We assume a fixed pretrained network for classification $f$ and a fixed pretrained generator $G$, for generation on the input dataset respectively. We use these two networks to guide our design and learning of $g$ and its concept extraction function $\Phi$. We first discuss modelling of $g$ by describing its constituents, $\Phi$ and $\Theta$.
@@
Viewability property . In order to improve visualization for unsupervised CoINs, and thus interpretation of learnt concepts, we propose to add the requirement for a viewability property. Given an input image $x$, this property requires to be able to reconstruct high-quality images from $\Phi(x)$. Specifically, reconstructions should have high enough quality to “view” input samples through generated outputs and thus ground modifications to $\Phi(x)$ back to $x$. We propose to achieve this by using a pretrained generative model $G$ and learning an additional concept translator module $\Omega$ to map $\Phi(x)$ to the latent space of $G$, such that high-quality reconstructed images can be obtained from $\Phi(x)$ through $\Omega$ and $G$. This also retains flexibility to design $\Phi$ for instance in choosing number of concepts $K$ according to the problem, regardless of the choice of pretrained $G$.
Pretrained generative model $G$ as decoder . In practice, we want our generative model to (i) have a low dimensional latent space, (ii) have a structured latent space that admits meaningful latent traversals, (iii) be able to generate high-quality images for the underlying data distribution. The choice of using a pretrained generator instead of simultaneously training is because it significantly lowers training costs, reduces training complexity and improves reusability. We discuss the significance of these properties in relation to various generative architectures in Appendix B. We also experimentally demonstrate the versatility of our method to different generative architectures in Appendix D.
Concept translator $\Omega$ Concept representations in CoINs are typically smaller dimensional than latent spaces of generative models as latent spaces encode lot more information about input than needed for classification. We thus learn the concept representation $\Phi(x)$ separately from the latent space of pretrained $G$ and instead learn a concept translator $\Omega$ to map $\Phi(x)$ to latent space of $G$. The design of $\Omega$ depends on the architecture of underlying generative model. In general, it consists of a fully-connected (FC) layer that predicts for each input image $x$, the latent vector $w_{x}$ from $\Phi(x)$, that will then be used as input for $G$. The computation of the reconstructed input $\tilde{x}$ is given by:
$\tilde{x}=G(w_{x}),\quad\text{where }w_{x}=\Omega(\Phi(x)).$ (1)
We discuss in more technical details, the architectures of each network in Appendix C.